OpenStudy (anonymous):

The perimeter of a rectangle can be found by using the formula P = 2l+2w. If the perimeter of rectangle BCND is 26 and BD = 5, find the length of the diagonal BN BN = 6.24 BN = 9.43 BN = 8 BN = 10

OpenStudy (anonymous):

26=2l+10 2l=16 l=8 \[\sqrt{8^2+5^2}\]BN=9.43